# Unavoidable patterns and plane paths in dense topological graphs

### v1

@@ 1 Introduction
A topological graph is a graph drawn in the plane such that its vertices are represented by points and its edges are represented by non-self-intersecting arcs connecting the corresponding points. No edge is allowed to pass through any point representing a vertex other than its endpoints. Tangencies between the edges are not allowed. That is, if two edges share an interior point, then they must properly cross at this point. A topological graph is called plane if there are no crossing edges. Given a topological graph $G$, we say that $H$ is a topological subgraph of $G$ if $V(H)\subset V(G)$ and $E(H)\subset E(G)$. We say that $G$ and $H$ are weakly isomorphic if there is an incidence preserving bijection between the vertices and edges of $G$ and $H$ such that two edges of $G$ cross if and only if the corresponding edges in $H$ cross as well. A topological graph is simple if every pair of its edges intersect at most once: at a common endpoint or at a proper crossing. Simple topological graphs are also known as a simple drawings. If the edges of a topological graph are drawn with straight-line segments, then it is called geometric. We call a geometric graph convex if its vertices are in convex position.
@@ 2 Unavoidable patterns in complete bipartite topological graphs
In this section, we prove Theorem 1.1. We will need the following combinatorial lemmas. Let $G$ be an ordered graph with vertex set $U=\{u_{1},u_{2},\ldots,u_{n}\}$. We say that $G$ contains a monotone path of length $k$ if there are $k+1$ indices $1\leq i_{0}<i_{1}<\cdots<i_{k}\leq$ such that $u_{i_{j}}u_{i_{j+1}}\in E(G)$ for all $0\leq i<k$. We say that $G$ is transitive if for any $1\leq i_{1}<i_{2}<i_{3}\leq n$, the condition $u_{i_{1}}u_{i_{2}},u_{i_{2}}u_{i_{3}}\in E(G)$ implies that $u_{i_{1}}u_{i_{3}}\in E(G)$. We will need the following three simple observations. The first is self evident, for the second we provide a simple proof to be self contained. The last one is a well known estimate of the multicolor Ramsey numbers.
@@
Let $r(k;r)$ be the minimum integer $n$ such that every $r$-coloring of the edges of the complete graph on $n$ vertices contains a monochromatic clique on $k$ vertices. The existence of $r(k;r)$ follows by the famous theorem of Ramsey [24], and following the arguments of Erdős-Szekeres [9], we have
@@
Figure 1: A drawing of each of the 55 types of $C_{4}$’s.
@@
We claim that the ordered graph $G^{(w)}_{s,t}$ is transitive for $w\in\{1,2,3,4\}$, that is, with respect to the first four order types. Indeed, it suffices to verify the statement for $G^{(1)}_{s,t}$, since a symmetric argument would follow for $G^{(2)}_{s,t},G^{(3)}_{s,t}$, and $G^{(4)}_{s,t}$. For $i<j<\ell$, suppose that $\{u_{i},u_{j}\}$ and $\{u_{j},u_{\ell}\}$ are edges in $G^{(1)}_{s,t}$. Consider the drawing of $K_{2,2}$ induced by the 4-tuple $\{u_{i},u_{j},v_{s},v_{t}\}$, specifically the “triangle” formed by the edge $u_{j}v_{t}$ and segments of the edges $u_{j}v_{s}$ and $u_{i}v_{t}$. Notice that $u_{\ell}$ cannot be inside this triangle otherwise $(u_{j},v_{s},v_{t},u_{\ell})$ appear in counter-clockwise order from the crossing point of $u_{j}v_{t}$ and $u_{\ell}v_{s}$. Since edge $u_{\ell}v_{s}$ must cross $u_{j}v_{t}$, it crosses one of the two other sides, but (as we have a simple drawing) it must not cross $u_{j}v_{s}$, so it crosses $u_{i}v_{t}$. Further, $u_{\ell}$, $v_{s}$, and $v_{t}$ must appear in the same cyclic order from the crossing point of $u_{j}v_{t}$ and $u_{\ell}v_{s}$ as well as from the crossing point of $u_{i}v_{t}$ and $u_{\ell}v_{s}$. The former is clockwise order from the assumption $\{u_{j},u_{\ell}\}\in E(G^{1})$, so the latter also clockwise implying $\{u_{i},u_{\ell}\}\in E(G_{1})$ as claimed. See Figure 2.
@@
This claim implies that we can label the edge $v_{s}v_{t}$ of the complete graph on $V$ with a pair $(S,i)$, where $i\in\{1,\ldots,5\}$ and $S\subset U$ is a clique in $G^{(i)}_{s,t}$ of size $3$ if $i=5$ or of size $k$ otherwise. In case there are more possible labels we choose arbitrarily. Since $|U|=2(k-1)^{4}+1\leq 2k^{4}$ and $k\geq 3$, the number of distinct labels we use on the edges is at most
$4\binom{|U|}{k}+\binom{|U|}{3}\leq k^{4k}.$
@@
Consider four vertices $u_{1},u_{2},u_{3},u_{4}\in U$ and that appear in this order on the $x$ axis. These vertices induce a drawing of $K_{4}$ in which only the edges $u_{1}u_{3}$ and $u_{2}u_{4}$ cross and only if they are drawn in the same half plane. If these two edges are drawn in distinct half planes we obtain a plane drawing of $K_{4}$.
@@
Now let $S$ and $T$ be two separated $k$-element set of vertices. We may assume that $S$ splits into $S_{1}$ and $S_{2}$ as above and $S_{1}$ is not empty (otherwise, switch the roles of $S$ and $T$). Let $v_{0}$ and $v_{1}$ be the leftmost and rightmost vertices in $T$, respectively, while $u_{1}$ is the rightmost vertex in $S_{1}$ and $u_{0}$ is the leftmost vertex in $S_{2}$, or in case $S_{2}$ is empty, then $u_{0}$ is the leftmost vertex in $S_{1}$. We claim that the subgraph formed by the edges between $S$ and $T$ is weakly isomorphic to $C_{k,k}$ if and only if all of its edges, with the possible exception of $u_{1}v_{0}$ and $u_{0}v_{1}$, are drawn in the same (upper or lower) half plane. The if part of this equivalence is easy to check (and strictly speaking, not even needed for our argument). For the only if part assume without loss of generality that $u_{1}v_{1}$ is drawn in the upper half plane. Now consider two other vertices: $u_{1}\neq u\in S$ and $v_{1}\neq v\in T$. If $uv$ drawn in the lower half plane, then the $4$-cycle $u_{1}v_{1}uv$ is a plane drawing, so we do not have the desired weak isomorphism. So the weak isomorphism implies that all of these edges are drawn in the upper half plane including the edge $u_{0}v_{0}$. But then a similar argument shows that all the edges $uv$ with $u_{0}\neq u\in S$ and $v_{0}\neq v\in T$ must also be drawn in the upper half plane. This covers all the edges mentioned in the claim.
@@ 3 Local thrackles
For the if part of proposition, notice that the components of $G$ can be drawn pairwise disjointly, so it is enough to prove that caterpillars, odd spiked cycles and even spiked cycles of length at least $8$ are geometric local thrackleable. The first two cases are already handled by Theorem 3.1 (as those graphs are even geometric thrackleable), so here we concentrate on even spiked cycles. Let $n\geq 8$ even and consider the vertices $v_{1},\dots,v_{n}$ of a regular $n$-gon. Choose m=n/21m=n/2-1 or n/22n/2-2 whichever is odd. Connect each $v_{i}$ to $v_{i+m}$ (indices understood modulo $n$). As $n$ and $m$ are relatively prime, this is a straight line drawing of $C_{m}$ and for $m\neq 10$ we have $3m>n$ making this a local thrackle. We can add an arbitrary number of “spikes” by connecting $v_{i}$ to vertices on the interval $v_{i+m}v_{i+m+1}$ and the geometric graph remains a local thrackle. See left side of Figure 4. The missing $n=10$ case is handled by the drawing on the right side of Figure 4.∎
Figure 4: Even cycles with spikes can be drawn as a geometric local thrackle.
@@
Delete at each left-vertex the incident edge that leaves highest at this vertex. Delete at each right-vertex the incident edge that leaves lowest at this vertex. We have deleted at most $l+r$ edges. Notice that if there is a remaining edge between a left-vertex and a right-vertex than together with the $2$ deleted edges incident to these vertices we get a plane $3$-path, a contradiction. Therefore, no edge is left between a left- and a right-vertex. Similarly, if a middle-vertex is connected to two left-vertices in the remaining graph, then adding the deleted edge of the left-vertex on the higher of these two edges, we obtain a plane $3$-path. The contradiction shows that every middle vertex is connected to at most one left-vertex in the remaining graph.
@@
Figure 6: The star drawn as a local thrackle.
@@
Finally, we are left to prove that this is a local thrackle. It is easy to check we a have simple drawing. It remains to check that no 3-edge path is planar. For this, notice that any $3$-path contains exactly one vertex from $V\cup W$. This implies that the $3$ edges of any $3$-path belong to a star that was drawn as a local thrackle. For $v_{1}$ this follows from our drawing of $S_{n}$, for other $v_{i}$ it follows from the fact that these edges go close to the respective edge from $v_{1}$, and for $w_{i}$ it again follows from the fact that at $w_{i}$ we drew the edges in distance two as a local thrackle. ∎
@@ 3.1 Further examples of local thrackles
In the proof of the previous claim, we have seen one operation that produces new local thrackleable graphs from existing ones. We show two more such operations.
@@
Let $G$ be a local thrackleable graph which has an induced $k$-path for some $k\geq 4$. Then if we add another copy of this path to $G$ with the same endvertices and new inner vertices to get the graph $G^{\prime}$, then $G^{\prime}$ is also thrackleable.
@@
Using the above claim and that all cycles are local thrackleable (except for $C_{4}$) we get that:
@@
Claim 3.12.
Let $G$ be a local thrackleable graph which has an induced $3$-path with middle edge $uv$. Then if we take two copies of $G$, $G_{1}$ and $G_{2}$ and add a new edge between $u_{2}$ and $v_{1}$ (the vertex corresponding to $u$ in $G_{2}$ and to $v$ in $G_{1}$), to get the graph $G^{\prime}$, then $G^{\prime}$ is also local thrackleable.
Proof.
We draw $G_{1}$ as a local thrackle and then draw $G_{2}$ so that the corresponding pairs of points are very close and the corresponding pairs of edges go very close and intersect each other at most once. Then we redraw the vertices $u_{2},v_{2}$ and $4$ incident edges as in Figure 10 (corresponding pairs of points are very close and corresponding pairs of edges go very close and intersect at most once). Note that we can draw $G_{2}$ such that the vertex neighboring $v_{2}$ is as in the figure, and then whatever is the position of the other endpoint of the $3$-path in $G_{2}$ (the one neighboring $u_{2}$), we can add the new edges so that the resulting graph is a local thrackle. See Figure 10. ∎
Figure 10: Duplicating a graph and adding one new edge.
This claim implies for example that two cycles of the same length of at least $4$ connected by an edge can be drawn as a local thrackle. We note that this claim can be easily generalized further to allow adding new edges at the same time at multiple vertex-disjoint induced $3$-paths.
Without going into details, we mention that with a combination of the $3$ operations described above, we can realize many more dumbbells and theta graphs as local thrackles.
@@ 5 Related problems and concluding remarks
2. The following 4-uniform hypergraph Ramsey problem was also studied independently by Negami [18] and Mathias Schacht. Let $S_{a,b}=(U,V,E)$ be the 4-uniform hypergraph with the vertex set the disjoint union of $U$ and $V$, where $|U|=a$ and $|V|=b$, such that $E(S_{a,b})=\{\{x,y,z,w\}:x\neq y,z\neq w,x,y\in U,\textnormal{ and }z,w\in V\}.$ What is the minimum integer $r(S_{n,n})=N$ such that for any red/blue coloring of the edges of $S_{N,N}$, there is a monochromatic copy of $S_{n,n}$? Following the arguments given in Section 2, one can show that $r(S_{n,n})<2^{2^{O(n^{2})}}$, while a standard probabilistic argument shows that $r(S_{n,n})>2^{\Omega(n^{3})}$.

### v2

@@ 1 Introduction
A topological graph is a graph drawn in the plane such that its vertices are represented by points and its edges are represented by non-self-intersecting arcs connecting the corresponding points. No edge is allowed to pass through any point representing a vertex other than its endpoints. Tangencies between the edges are not allowed. That is, if two edges share an interior point, then they must properly cross at this point. A topological graph is called plane if there are no crossing edges. Given a topological graph $G$, we say that $H$ is a topological subgraph of $G$ if $V(H)\subset V(G)$ and $E(H)\subset E(G)$. We say that $G$ and $H$ are weakly isomorphic if there is an incidence preserving bijection between the vertices and edges of $G$ and $H$ such that two edges of $G$ cross if and only if the corresponding edges in $H$ cross as well. A topological graph is simple if every pair of its edges intersect at most once: at a common endpoint or at a proper crossing. Simple topological graphs are also known as simple drawings. If the edges of a topological graph are drawn with straight-line segments, then it is called geometric. We call a geometric graph convex if its vertices are in convex position.
@@ 2 Unavoidable patterns in complete bipartite topological graphs
In this section, we prove Theorem 1.1. We will need the following combinatorial lemmas. Let $G$ be an ordered graph with vertex set $U=\{u_{1},u_{2},\ldots,u_{n}\}$. We say that $G$ contains a monotone path of length $k$ if there are $k+1$ indices $1\leq i_{0}<i_{1}<\cdots<i_{k}\leq n$ such that $u_{i_{j}}u_{i_{j+1}}\in E(G)$ for all $0\leq i<k$. We say that $G$ is transitive if for any $1\leq i_{1}<i_{2}<i_{3}\leq n$, the condition $u_{i_{1}}u_{i_{2}},u_{i_{2}}u_{i_{3}}\in E(G)$ implies that $u_{i_{1}}u_{i_{3}}\in E(G)$. We will need the following three simple observations. The first is self evident, for the second we provide a simple proof to be self contained. The last one is a well known estimate of the multicolor Ramsey numbers.
@@
Let $r(k;r)$ be the minimum integer $n$ such that every $r$-coloring of the edges of the complete graph on $n$ vertices contains a monochromatic clique on $k$ vertices. The existence of $r(k;r)$ follows by the famous theorem of Ramsey [24], and following the arguments of Erdős–Szekeres [9], we have
@@
Figure 1: A drawing of each of the five types of $C_{4}$’s.
@@
We claim that the ordered graph $G^{(w)}_{s,t}$ is transitive for $w\in\{1,2,3,4\}$, that is, with respect to the first four order types. We verify the statement for $G^{(1)}_{s,t}$. For $i<j<\ell$, suppose that $\{u_{i},u_{j}\}$ and $\{u_{j},u_{\ell}\}$ are edges in $G^{(1)}_{s,t}$. Consider the drawing of $K_{2,2}$ induced by the 4-tuple $\{u_{i},u_{j},v_{s},v_{t}\}$, specifically the “triangle” formed by $u_{j}$, $v_{t}$ and the crossing point between $u_{j}v_{s}$ and $u_{i}v_{t}$. Since $(u_{j},v_{s},v_{t},u_{\ell})$ must appear in clockwise order from the crossing point of $u_{j}v_{t}$ and $u_{\ell}v_{s}$, we must exit the “triangle” at this crossing point as we travel along $u_{\ell}v_{s}$ from $v_{s}$ to $u_{\ell}$. This implies that $u_{\ell}v_{s}$ crosses both $u_{j}v_{t}$ and $u_{i}v_{t}$. Further, $u_{\ell}$, $v_{s}$, and $v_{t}$ must appear in the same cyclic order from the crossing point of $u_{j}v_{t}$ and $u_{\ell}v_{s}$ as well as from the crossing point of $u_{i}v_{t}$ and $u_{\ell}v_{s}$. The former is clockwise order from the assumption $\{u_{j},u_{\ell}\}\in E(G^{(1)}_{s,t})$, so the latter also clockwise implying $\{u_{i},u_{\ell}\}\in E(G^{(1)}_{s,t})$ as claimed. See Figure 2. The transitivity property of $G^{(2)}_{s,t},G^{(3)}_{s,t}$, and $G^{(4)}_{s,t}$ can be argued similarly, or be reduced to the case of $G^{(1)}_{s,t}$ by reversing the vertex orders on $U$ or $V$. For example, a type 3 quadruple $\{u_{i},u_{j},v_{s},v_{t}\}$ would have been type 1 if we had ordered the vertices in $U$ the same but in $V$ reversely.
@@
This claim implies that we can label the edge $v_{s}v_{t}$ of the complete graph on $V$ with a pair $(S,i)$, where $i\in\{1,\ldots,5\}$ and $S\subset U$ is a clique in $G^{(i)}_{s,t}$ of size $3$ if $i=5$ or of size $k$ otherwise. In case there are more possible labels we choose arbitrarily. For $k\geq 7$, using $|U|=2(k-1)^{4}+1\leq 2k^{4}$, the number of distinct labels we use on the edges is at most
$r(k):=4\binom{|U|}{k}+\binom{|U|}{3}\leq 5\binom{2k^{4}}{k}\leq 5(2ek^{3})^{k}\leq k^{4k}.$
And we can further estimate using $k\geq 7$
$r(k)^{r(k)\cdot k}<2^{k^{5k}}.$
For $k=3,4,5,6$, the upper bound for $r(k)$ is too crude, but we can verify the previous inequality by direct computation.
@@
Consider four vertices $u_{1},u_{2},u_{3},u_{4}\in U$ that appear in this order on the $x$ axis. These vertices induce a drawing of $K_{4}$ in which only the edges $u_{1}u_{3}$ and $u_{2}u_{4}$ cross if and only if they are drawn in the same half plane. If these two edges are drawn in distinct half planes we obtain a plane drawing of $K_{4}$.
@@
Now let $S$ and $T$ be two separated $k$-element set of vertices. We may assume that $S$ splits into $S_{1}$ and $S_{2}$ as above and $S_{1}$ is not empty (otherwise, switch the roles of $S$ and $T$). Let $v_{0}$ and $v_{1}$ be the leftmost and rightmost vertices in $T$, respectively, while $u_{1}$ is the rightmost vertex in $S_{1}$ and $u_{0}$ is the leftmost vertex in $S_{2}$, or in case $S_{2}$ is empty, then $u_{0}$ is the leftmost vertex in $S_{1}$. We claim that the subgraph formed by the edges between $S$ and $T$ is weakly isomorphic to $C_{k,k}$ if and only if all of its edges, with the possible exception of $u_{1}v_{0}$ and $u_{0}v_{1}$, are drawn in the same (upper or lower) half plane. The if part of this equivalence is easy to check (and strictly speaking, not even needed for our argument). For the only if part assume without loss of generality that $u_{1}v_{1}$ is drawn in the upper half plane. Now consider two other vertices: $u_{1}\neq u\in S$ and $v_{1}\neq v\in T$. If $uv$ is drawn in the lower half plane, then the $4$-cycle $u_{1}v_{1}uv$ is a plane drawing, so we do not have the desired weak isomorphism. So the weak isomorphism implies that all of these edges are drawn in the upper half plane including the edge $u_{0}v_{0}$. But then a similar argument shows that all the edges $uv$ with $u_{0}\neq u\in S$ and $v_{0}\neq v\in T$ must also be drawn in the upper half plane. This covers all the edges mentioned in the claim.
@@ 3 Local thrackles
For the if part of proposition, notice that the components of $G$ can be drawn pairwise disjointly, so it is enough to prove that caterpillars, odd spiked cycles and even spiked cycles of length at least $8$ are geometric local thrackleable. The first two cases are already handled by Theorem 3.1 (as those graphs are geometric thrackleable), so here we concentrate on even spiked cycles. Let $n\geq 8$ be even. We first place vertices $v_{1},v_{2},\dots,v_{n-3}$ on the circle according to the following clockwise order: $v_{1},v_{3},v_{5},\dots,v_{n-3}$, $v_{2},v_{4},\dots,v_{n-4}$. Then we put $v_{n-2}$ on the clockwise arc $v_{1}v_{3}$, $v_{n-1}$ on the clockwise arc $v_{2}v_{4}$, and $v_{n}$ on the clockwise arc $v_{n-5}v_{n-3}$. Observe that we obtain a geometric local threackle drawing of $C_{n}$ by connecting $v_{i}v_{i+1}$ using straight-line segments. See Figure 4 for an illustration. Finally, we can choose an interval on the cycle for each $v_{i}$ such that if we place the leaf neighbors of $v_{i}$ on this intervals the resulting drawing is still a local thrackle. ∎
Figure 4: $C_{8}$ can be drawn as a linear local thrackle, and with a spike $u$ attached to $v_{1}$.
@@
At each left-vertex, delete the incident edge that leaves highest at this vertex. Delete at each right-vertex the incident edge that leaves lowest at this vertex. We have deleted at most $l+r$ edges. Notice that if there is a remaining edge between a left-vertex and a right-vertex, then together with the $2$ deleted edges incident to these vertices, we get a plane $3$-path, a contradiction. Therefore, no edge is left between a left- and a right-vertex. Similarly, if a middle-vertex is connected to two left-vertices in the remaining graph, then adding the deleted edge of the left-vertex on the higher of these two edges, we obtain a plane $3$-path. The contradiction shows that every middle vertex is connected to at most one left-vertex in the remaining graph.
@@
Figure 6: The 1-subdivision of a star drawn as a local thrackle.
@@
Finally, we are left to prove that this is a local thrackle. It is easy to check we have a simple drawing. It remains to check that no 3-edge path is planar. For this, notice that any $3$-path contains exactly one vertex from $V$ and $W$ respectively. This implies that the $3$ edges of any $3$-path belong to a 1-subdivision of a star that was drawn as a local thrackle. For $v_{1}$ this follows from our drawing of $S_{n}$, for other $v_{i}$ it follows from the fact that these edges go close to the respective edge from $v_{1}$, and for $w_{i}$ it again follows from the fact that at $w_{i}$ we drew the edges in distance two as a local thrackle. ∎
@@ 3.1 Further examples of local thrackles
In the proof of the previous claim, we have seen one operation that produces new local thrackleable graphs from existing ones. We show one more such operation.
@@
Let $G$ be a local thrackleable graph which has an induced $k$-path for some $k\geq 4$. Then if we add another copy of this path to $G$ with the same endvertices and new inner vertices to get the graph $G^{\prime}$, then $G^{\prime}$ is also local thrackleable.
@@
Using the above claim and that all cycles are local thrackleable (except for $C_{4}$) among others we get that:
@@ 5 Related problems and concluding remarks
2. The following 4-uniform hypergraph Ramsey problem was also studied independently by Negami [18] and Mathias Schacht (private communication). Let $S_{a,b}=(U,V,E)$ be the 4-uniform hypergraph with the vertex set the disjoint union of $U$ and $V$, where $|U|=a$ and $|V|=b$, such that $E(S_{a,b})=\{\{x,y,z,w\}:x\neq y,z\neq w,x,y\in U,\textnormal{ and }z,w\in V\}.$ What is the minimum integer $r(S_{n,n})=N$ such that for any red/blue coloring of the edges of $S_{N,N}$, there is a monochromatic copy of $S_{n,n}$? Following the arguments given in Section 2, one can show that $r(S_{n,n})<2^{2^{O(n^{2})}}$, while a standard probabilistic argument shows that $r(S_{n,n})>2^{\Omega(n^{3})}$.